# ChainBuddy: An AI Agent System for Generating LLM Pipelines

### v1

@@ Abstract.
As large language models (LLMs) advance, their potential applications have grown significantly. However, it remains difficult to evaluate LLM behavior on user-specific tasks and craft effective pipelines to do so. Many users struggle with where to start, often referred to as the ”blank page problem.” ChainBuddy, an AI assistant for generating evaluative LLM pipelines built into the ChainForge platform, aims to tackle this issue. ChainBuddy offers a straightforward and user-friendly way to plan and evaluate LLM behavior, making the process less daunting and more accessible across a wide range of possible tasks and use cases. We report a within-subjects user study comparing ChainBuddy to the baseline interface. We find that when using AI assistance, participants reported a less demanding workload and felt more confident setting up evaluation pipelines of LLM behavior. We derive insights for the future of interfaces that assist users in the open-ended evaluation of AI.
@@
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Pre-print; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Despite these advancements, many users encounter a significant challenge: the ”blank page problem.” This problem is characterized by the uncertainty and difficulty of knowing where to begin when using platforms like promptfoo (Webster, 2023) or Flowise (Flowise AI, 2024). Some researchers have proposed social solutions to this problem, such as the development of prompt repositories and community-sharing mechanisms (Zamfirescu-Pereira et al., 2023). Another proposed solution is to create an assistant that helps users get started by providing initial guidance and structure (Arawjo et al., 2024).
Taking the latter approach, our solution, ChainBuddy, is an AI-powered assistant that automatically generates starter LLM pipelines (“flows”) given an initial prompt. ChainBuddy helps users get started with evaluating LLM behavior and setting up chains by providing a starter flow, custom-tailored to their use case, that they can edit and extend. We built ChainBuddy on top of an existing open-source visual environment, ChainForge (Arawjo et al., 2024), which was designed for open-ended prompt engineering, LLM evaluation, and experimentation tasks (comparing response quality across prompts and models, setting up data processing pipelines, and establish automated evaluation metrics). By building on top of this foundation, ChainBuddy offers users structured assistance to help them overcome the blank page problem, making it easier to explore, experiment and evaluate the behavior of LLMs across a wide range of potential tasks and use cases. We report a mixed methods usability study of ChainBuddy, comparing the assistant to the baseline interface. We find broad support for ChainBuddy, particularly in its reduction of user effort and requirements-gathering intent elicitation feature, with the majority of participants expressing surprise at the quality of the assistant’s capabilities. Our contributions are:
@@ 2. Related Work
The problem of pipeline generation bears a similarity to AutoML, an area of research in machine learning that focuses on automatically designing machine learning (ML) pipelines to train new ML models, whether in part or in full (Heffetz et al., 2020; Barbudo et al., 2023; Feurer et al., 2022). Certainly, AutoML faces unique challenges compared to the typical problems in the LLMOps space (such as managing very large training datasets, deciding upon hyperparameters and weighing trade-offs in terms of expected training cost and performance). Inspired by this line of ML research, here we investigate the end-to-end generation of LLM pipelines, an emerging research area we call AutoLLMOps: from a single user prompt, can we generate an inspectable, interactive, and editable pipeline, complete with input data, prompt(s) and model(s), and even automated evaluations? Could we build such a system for open-ended tasks? What benefits would users derive from it? And what dangers are there, if any, to automation? To the best of our knowledge, no HCI paper has yet investigated this question. With new LLM agent frameworks, emerging research is showing that this kind of end-to-end generation of a workflow is possible, albeit for different tasks like descriptive analytics given a user-provided dataset (Beasley and Abouzied, 2024) or for chatbot creation (Sánchez Cuadrado et al., 2024). Here, we apply this idea to LLM pipelines, specifically focusing on pipelines that help users evaluate LLM behavior.
@@ 3.2. System Architecture
Figure 3. ChainBuddy system architecture. A front-end requirement agent elicits user intent and context (left). When the user presses the generate button (Fig 2d), a specification of user intent is sent to the Planner in the back-end. The Planner agent breaks down the problem into tasks and sends each to a dedicated agent; the outputs are combined through layout-providing connection agents, merged and passed to an optional Reviewer agent. The final output is passed to the front-end as a complete ChainForge flow (JSON).
@@ 3.2.2. Workflow generation
  • Planner Agent: Takes the specification from the front-end of the user goal, and develops a comprehensive plan for implementation. The Planner is passed contextual information on all nodes in ChainForge that it has access to, their names and descriptions, and how they are allowed to connect.
@@
  • Post-hoc Reviewer Agent: A final Reviewer agent assesses the generated flow against the initial user criteria given to the Planner agent. The Reviewer can trigger the Planner to re-plan before presenting the flow to the user, if the generated flow falls below expectations. (Note that for our usability study, we disabled the Reviewer because it significantly increases generation time.)
By leveraging these architectural principles, ChainBuddy aims to provide a robust, scalable solution for a wide range of LLM-related tasks. Note finally that the system we present here is limited to a few select nodes in ChainForge: TextFields Node, for defining input data; Prompt Node, for prompting one or more models and prompt templating; Python Code Evaluator and LLM Scorer Nodes for evaluating LLM outputs; and the Vis Node. The system also supports template chaining (Arawjo et al., 2024) to compare across prompts; i.e., putting prompt templates inside TextField inputs and chaining them together. Template chaining is primarily useful for comparing across prompts in a structured way.
@@ 3.3. Early Feedback
  • Structured elicitation: We implemented ChatGPT-like chat where the system asks users questions. However, the LLM output was often long or listed several questions, and users struggled to reply in a natural way. Based on this, we opted for a structured form-filling approach, letting the user clarify only the questions they want to. We also limited the number and length of questions.
@@ 4. Usability Study
We also sought quantitative, subjective metrics for the following hypotheses for the structured tasks measured via self-reported completion times, [^2] Specifically, the time from when participants first clicked or typed in the interface, to the time when they said they were “done” or ran out of time. Participants were given 12 minutes for all tasks; if they ran out of time, the time was recorded as the maximum. NASA TLX cognitive load scale (Hart, 1988), and a subset of five system usability Likert questions [^3] The five questions, “I found the system easy to use,” “I would like to use this system frequently,” “I found the system unnecessarily complex,” “I felt confident using the system”, and “I needed to learn a lot of things before I could get going with this system”. derived from the System Usability Scale (Brooke, 1996):
@@ 4.2. Methodology, Procedure and Tasks
    A tutorial guides users through the process of writing a prompt template that takes an ingredient and returns the dishes it can make, comparing prompts using template chaining, querying models, and inspecting results. The tutorial gave participants everything they needed to succeed, as the solutions to tasks were structurally similar to the tutorial flow. (The Tutorial did not introduce the Assistant; see below.)
@@ 5. Findings
Figure 4. Participant responses to Likert Questions for NASA TLX and system usability, grouped by Condition. Asterisk (*) indicates a significant main effect of Condition at $p<0.05$; † indicates another main effect, interaction effect or outliers.
@@ 5.1.4. Performance (H4)
We estimated user performance on the task with three binary (true-false) measures shared across both task descriptions: whether their final flow explicitly compares two prompts, whether the flow runs (i.e., the Prompt Node can run), and whether the flow uses template chaining to compare prompts (the last is a feature introduced in the tutorial and was the expected solution, but is not the only one). Plots show strong evidence in favor of the Assistant condition (Figure 5), indicating that, despite participants rating themselves of having similar success in both interfaces, two-thirds of participants did not adequately solve the task in the Control condition (did not create a workflow helping users to compare prompts), compared to eleven (out of 12) participants when using the Assistant.
@@ 5.1.5. Number and Types of Created Nodes
Figure 5. Performance metrics comparing expected qualities of correct solutions across conditions and tasks: whether the flow explicitly compares two prompts, whether the flow runs (i.e., the Prompt Node can run), and whether the flow uses template chaining to compare prompts (a feature introduced in the tutorial, and the expected solution).

### v2

@@ Abstract.
As large language models (LLMs) advance, their potential applications have grown significantly. However, it remains difficult to evaluate LLM behavior on user-defined tasks and craft effective pipelines to do so. Many users struggle with where to start, often referred to as the ”blank page problem.” ChainBuddy, an AI workflow generation assistant built into the ChainForge platform, aims to tackle this issue. From a single prompt or chat, ChainBuddy generates a starter evaluative LLM pipeline in ChainForge aligned to the user’s requirements. ChainBuddy offers a straightforward and user-friendly way to plan and evaluate LLM behavior and make the process less daunting and more accessible across a wide range of possible tasks and use cases. We report a within-subjects user study comparing ChainBuddy to the baseline interface. We find that when using AI assistance, participants reported a less demanding workload, felt more confident, and produced higher quality pipelines evaluating LLM behavior. However, we also uncover a mismatch between subjective and objective ratings of performance: participants rated their successfulness similarly across conditions, while independent experts rated participant workflows significantly higher with AI assistance. Drawing connections to the Dunning–Kruger effect, we draw design implications for the future of workflow generation assistants to mitigate the risk of over-reliance.
@@
journalyear: 2025copyright: noneconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714085isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCI
[…]
ChainBuddy workflow for analyzing LLM responses based on persona-driven prompts. This interface captures a workflow generated by ChainBuddy, an assistant chat window in the bottom-right where the goal is to assess how different personas influence the responses of various LLMs to a complex math problem. The screen displays various UI elements such as ’Education Personas’, ’Science Personas’, and ’Humanities Personas’ panels, each containing different personas like ’mathematician’, ’physicist’, or ’poet’. A ’Math Question’ node featuring the Gaussian integral task connects to a ’Prompt Node’. The Prompt Node receives input from persona and question nodes, configuring prompts that tailor the question to each persona. This node feeds into a Python Code Evaluator script designed to validate LLM outputs against the correct answer, depicted by regex patterns for mathematical notation. The interface also includes response inspection options for comparing results across four different LLM versions: GPT-3.5, GPT-4, Gemini, and Claude. A visualization node presents a bar graph summarizing the number of ’true’ evaluations per persona and LLM model.
@@ 1. Introduction
Despite these advancements, many users encounter a significant challenge: the ”blank page problem.” This problem is characterized by the uncertainty and difficulty of knowing where to begin when using platforms like promptfoo (Webster, 2023) or Flowise (Flowise AI, 2024). Some researchers have proposed social solutions to this problem, such as the development of prompt repositories and community-sharing mechanisms (Zamfirescu-Pereira et al., 2023). Another proposed solution is to create an assistant that helps users get started by providing initial guidance and structure (Arawjo et al., 2024; Gordon et al., 2023).
Taking the latter approach, our solution, ChainBuddy, is an AI-powered workflow generation assistant that automatically generates starter LLM pipelines (“flows”) given an initial prompt. ChainBuddy helps users get started with evaluating LLM behavior and setting up chains by providing a starter flow, custom-tailored to their use case, that they can edit and extend. We built ChainBuddy on top of an existing open-source visual environment, ChainForge (Arawjo et al., 2024), which was designed for open-ended prompt engineering, LLM evaluation, and experimentation tasks (comparing response quality across prompts and models, setting up data processing pipelines, and establish automated evaluation metrics). By building on top of this foundation, ChainBuddy offers users structured assistance to help them overcome the blank page problem, making it easier to explore, experiment and evaluate the behavior of LLMs across a wide range of use cases. We report a mixed methods usability study of ChainBuddy, comparing the assistant to the baseline interface. We find broad support for ChainBuddy, particularly in its reduction of user effort and requirements-gathering intent elicitation feature, with the majority of participants expressing surprise at the quality of the assistant’s capabilities. Our contributions are:
@@
In particular, our user study reveals two complications for workflow generation assistants: that participants did not appear able to adequately assess the quality of their own work compared to the opinions of expert raters (see Sections 5.1.4 and 6); and that while workflow generation assistants can support onboarding, they do so at the potential risk of user over-reliance on AI outputs. In our Discussion, we reflect on these issues and suggest future work explores techniques and study designs that attend to them.
@@ 2. Related Work
The problem of pipeline generation bears a similarity to AutoML, an area of research in machine learning that focuses on automatically designing machine learning (ML) pipelines to train new ML models, whether in part or in full (Heffetz et al., 2020; Barbudo et al., 2023; Feurer et al., 2022). Certainly, AutoML faces unique challenges compared to the typical problems in the LLMOps space (such as managing very large training datasets, deciding upon hyperparameters and weighing trade-offs in terms of expected training cost and performance). Inspired by this line of ML research, here we investigate the end-to-end generation of LLM pipelines, an emerging research area we call AutoLLMOps: from a single user prompt, can we generate an inspectable, interactive, and editable pipeline, complete with input data, prompt(s) and model(s), and even automated evaluations? Could we build such a system for open-ended tasks? What benefits would users derive from it? And what dangers are there, if any, to automation? This problem bears a similarity to previous work on crowdsourcing-based task decomposition (Kittur et al., 2011; Kulkarni et al., 2012). With new LLM agent frameworks, emerging research is showing that this kind of end-to-end generation of a workflow is possible for tasks like descriptive analytics given a user-provided dataset (Beasley and Abouzied, 2024), visual programming (Zhou et al., 2023), AI agent workflows (Cheng et al., 2024), or for chatbot creation (Sánchez Cuadrado et al., 2024). Here, we specifically focus on generating LLM pipelines that help users evaluate LLM behavior.
@@ 3.1. Interface and Example Usage
[…]
ChainBuddy Interface and Example Usage. This figure demonstrates a step-by-step interaction between a user and the ChainBuddy system for generating tweet content from long text paragraphs. (A) A user initiates the interaction by specifying their need for tweet generation from long paragraphs. (B) ChainBuddy responds by presenting a structured form to the user, asking about preferences such as tweet aspects (brevity, humor), number of prompts, and specific constraints like character limits. (C) The user fills out and submits the form or engages in an open-ended chat to refine their input. (D) The user then presses a green button to generate the workflow. (E) After a brief processing delay, ChainBuddy outputs a starter pipeline. This pipeline includes long text paragraphs, a selection of prompt templates for tweet generation, a choice of LLMs to query (GPT-3.5 and GPT-4), and a Python-based code evaluator designed to ensure the responses adhere to the specified character limit.
@@ 3.2. System Architecture
[…]
ChainBuddy System Architecture. This diagram illustrates the structured workflow of ChainBuddy, which begins with a user interacting with a front-end requirement-gathering interface (left side). Here, a Requirement Agent captures user inputs and updates a chain of goal reasoning components, which includes the user’s primary goal, a list of requirements, and user preferences. Upon user confirmation (akin to pressing the generate button shown in Figure 2D), this refined user intent is sent to a Planner located in the back-end. The Planner decomposes the user’s problem into specific tasks and assigns them to specialized task-specific agents, such as text fields, prompts, and evaluators. Each task agent processes its part and passes the results to connection agents (Edge Agent and Position Agent), which structure the workflow layout.The final stage involves merging these components into a comprehensive JSON output, representing a complete ChainForge flow, which is sent back to the user’s front-end for interaction or further modification.
Figure 3. ChainBuddy system architecture. A front-end requirement agent elicits user intent and context (left). When the user presses the generate button (Fig 2d), a specification of user intent is sent to the Planner in the back-end. The Planner agent breaks down the problem into tasks and sends each to a dedicated agent; the outputs are combined through layout-providing connection agents and merged. The final output is passed to the front-end as a complete ChainForge flow (JSON).
@@ 3.2.2. Workflow generation
  • Planner Agent: Takes the specification from the front-end of the user goal, and develops a comprehensive plan for implementation. The Planner is passed contextual information on all nodes in ChainForge that it has access to, their names and descriptions, and how they are allowed to connect. For full details, please see Appendix B.
@@
A detailed walkthrough of how our system goes from initial user requirements to the final workflow, alongside example prompt templates and outputs from individual agents, is presented in Appendix B. Note that the system we present here is limited to a few select nodes in ChainForge: TextFields Node, for defining input data; Prompt Node, for prompting one or more models and prompt templating; Python Code Evaluator and LLM Scorer Nodes for evaluating LLM outputs; and the Vis Node. The system also supports template chaining (Arawjo et al., 2024) to compare across prompts; i.e., putting prompt templates inside TextField inputs and chaining them together. Template chaining is primarily useful for comparing across prompts in a structured way. In Appendix B, we cover how consistency in variable names is accomplished via variable traces, which enables template chaining.
@@ 3.3. Early Feedback
  • Structured elicitation: We implemented ChatGPT-like chat where the system asks users questions. We first tried in ‘infinite questioner’ approach where the agent keeps asking the user disambiguating questions, akin to Cheng et al. (Cheng et al., 2024); however, we found this approach slow and stressful to the user, as the user feels pressured to reply to every question, or struggled to reply in a natural way. The LLM output was also often long or listed several questions at once. Based on this feedback, we opted for a structured form-filling approach, presenting three questions at once and letting the user address only the questions they wanted to. We also limited the number and length of questions.
@@
3.4. Technical Evaluation
Following our iterative design process, we conducted a small-scale technical evaluation to estimate the quality of ChainBuddy’s pipeline generation algorithm. Evaluating workflows is an active area of research and is not straightforward: in many situations, as we reflect on in Discussion, there is no one ground truth “best” workflow; what is more, there is no benchmark of evaluative LLM pipelines, and our prototype was deliberately limited its featureset (Section 3.2.2). We thus prepared a set of representative prompts to the system. We used the same three broad categories of use cases found in a prior study of the base interface (Arawjo et al., 2024)—Prompt Engineering, LLM Auditing, and Data Processing—with corresponding subtasks (e.g., “comparing across prompts”) and three prompts per subtask, for a total of 27 prompts (Appendix A). These prompts were devised such that with its current features, the system should reasonably be able to address them (e.g., we did not include ground truth evaluation with a spreadsheet, as the Tabular Data node was not part of our prototype), but without foreknowledge of how the system would perform. Additionally, the advising author, and not the author who programmed the system, wrote the prompts. Each prompt was provided to ChainBuddy verbatim (no requirements gathering step). The generated workflow was run and exported, with only the layout adjusted for ease of analysis (e.g., moving nodes around, as our layout system is not perfect).
To grade workflow quality, we recruited 6 expert raters who are AI/ML experts with experience prompt engineering, [^2] In the screening survey, all participants indicated they have “Advanced (extensive experience)” or “Intermediate (regularly work with LLMs)” levels of experience with both LLMs and prompt engineering. In addition, when asked their level of Python programming skill, four wrote a 5 and two a 4 (out of 5), indicating advanced knowledge; this was important for evaluating Code Evaluator nodes. Participants were recruited from an institute for AI research. and paid them \$75 CAD each for 75 minutes of time. Each rater was assigned one prompt from each subtask, for a total of 9 workflows, with the exact prompts assigned at random in a balanced manner such that two expert raters independently graded each workflow. Raters were asked to grade how well each workflow aligned with the input requirements on a scale of 1-5 (Appendix A.1), and to provide qualitative feedback and reasoning for their score. The total 27 prompts are listed in Appendix A.2, alongside quality scores averaged across two independent raters, and a summary of qualitative feedback.
Across all categories, the average rating is 4.09 with a standard deviation of 0.7, indicating the vast majority of ratings are within the 3-5 range. A score of 4 indicates that the average workflow generated by ChainBuddy was aligned to the user requirements, with only 1-2 minor issues that were easily fixable (Appendix A.1). Notably, only one workflow obtained an average score less than 3 (prompt Q3.3.1). Running a one-way ANOVA on the average scores, comparing across categories, is not significant ($p{=}0.42$); thus, ChainBuddy had similar performance across categories considered here. 22 out of 27 ratings were within one point of each other, with raters expressing slight agreement (weighted Cohen’s kappa = 0.17).
Overall, this indicated that ChainBuddy was strong enough to proceed to a user study; yet also indicated areas for improvement. From the qualitative feedback, the most consistent issue across 14 workflows concerned the quality of the evaluator(s) implemented and chosen: LLM Scorers could have prompts with criteria that were slightly too vague; and Code Evaluators could be “too stringent” and struggle to generalize (e.g., to different formatting of LLM outputs). Beyond this, 6 issues were related to the Vis Node plot, specifically not plotting by a certain expected variable like prompt or prompt variable. This was caused by an oversight on our part: our ChainBuddy prototype only added a default Vis Node which by default plots by LLM; while the broader ChainForge interface allows users to change the variable. Although we also saw minor problems with input data and phrasing of prompts, this feedback indicates that most challenges stem from choices ChainBuddy makes in the evaluation part of LLM pipelines. Note that aligning evaluators with user expectations is non-trivial, with one recent solution arguing the necessity of a human-in-the-loop approach to align generated evaluators through user grading of LLM outputs (Shankar et al., 2024b).
@@ 4. Usability Study
We also sought quantitative, subjective metrics for the following hypotheses for the structured tasks measured via self-reported completion times, [^3] Specifically, the time from when participants first clicked or typed in the interface, to the time when they said they were “done” or ran out of time. Participants were given 12 minutes for all tasks; if they ran out of time, the time was recorded as the maximum. NASA TLX cognitive load scale (Hart and Staveland, 1988), and a subset of five system usability Likert questions [^4] The five questions, “I found the system easy to use,” “I would like to use this system frequently,” “I found the system unnecessarily complex,” “I felt confident using the system”, and “I needed to learn a lot of things before I could get going with this system”. derived from the System Usability Scale (Brooke, 1995):
@@ 4.2. Methodology, Procedure and Tasks
    A tutorial that guides users through the process of writing a prompt template that takes an ingredient and returns the dishes it can make, comparing prompts using template chaining, querying models, and inspecting results. The tutorial gave participants everything they needed to succeed, as the solutions to tasks were structurally similar to the tutorial flow. (The Tutorial did not introduce the Assistant; see below.)
@@ 5. Findings
[…]
The figure consists of a series of side-by-side boxplots comparing the results of Likert scale ratings for two conditions, labeled “Assistant” (red) and “Control” (blue). Each boxplot corresponds to a different attribute rated on a 5-point Likert scale, with 1 representing the lowest score and 5 representing the highest. The attributes evaluated are: ’Confident’, ’EasyToUse’, ’HardToLearn’, ’HardWork’, ’LikeToUse’, ’Mental’, ’Physical’, ’Rushed’, ’Stressed’, ’Successful’, ’TooComplex’ Each boxplot displays the distribution of responses for the Assistant and Control conditions. The red and blue boxes represent the interquartile range (IQR) between the first and third quartiles, with the line inside each box indicating the median value. The whiskers extend to the minimum and maximum values that are not considered outliers, and any points beyond the whiskers represent potential outliers. Below the boxplots is a legend, where the red box corresponds to “Assistant” and the blue box to “Control.”An asterisk (*) next to a condition name denotes a significant main effect where the Assistant condition differs from the Control at p ¡ 0.05. A cross (†) indicates other significant effects or outliers, highlighting more complex interactions or exceptions in the data. Metrics marked with an asterisk include ’Confident’, ’Mental’, and ’Physical’, indicating significantly higher scores in these aspects when using the Assistant. A cross (†) indicating other significant effects or outliers appears next to ’EasyToUse’, ’Successful’.
Figure 4. Participant responses to Likert Questions for six NASA TLX measures and five system usability measures, grouped by Condition. Asterisk (*) indicates a significant main effect of Condition at $p<0.05$; † indicates another main effect, interaction effect or outliers.
@@ 5.1.4. Performance (H4)
To examine performance, we asked the same expert raters from Section 3.4 to rate the workflows of participants on a scale from 1-5 (Appendix A.1). Experts were blinded to condition, given the same task description as participants, and randomly assigned 8 workflows, with each expert receiving an equal number of workflows from each condition (4 each). For reliability, two separate experts rated each user workflow.
A Mann-Whitney U test shows significant gains in performance for the AI assisted condition compared to the Control at $p{<}0.001$ ($U{=}108$; see Figure 5). Qualitative examination of the workflows suggests that two-thirds of the workflows in the human condition failed to compare two prompts, compared to all but one workflow in the AI condition; and two-thirds of workflows did not use template chaining, which was an expected part of the user’s solution (but not the only valid method). A weighted Cohen’s kappa provides an estimate of inter-rater agreement [^5] Note that because of the particular setup of random assignment to raters, this is only an estimate; we provide it to help readers parse our results. For our MWU test reported here, we included all datapoints; but note that averaging the two rater scores for each workflow and conducting MWU also reaches significance at $p=0.001$. at $k{=}0.388$, indicating fair agreement ($z{=}1.96$, $p{=}0.0499$) among raters.
@@ 5.1.5. Number and Types of Created Nodes
[…]
This bar chart compares expert ratings of user performance in two conditions: ”With AI Support” (top bar) and ”Without AI Support” (bottom bar). Each bar represents the proportion of participants receiving grades on a scale from 1 (lowest) to 5 (highest), with darker shades of green indicating higher grades. In the ”With AI Support” condition, a greater proportion of participants received higher grades (4 and 5), while the ”Without AI Support” condition shows more participants receiving lower grades (1 and 2). The chart emphasizes the improvement in performance with AI support.
Figure 5. Performance ratings of participant workflows by expert raters, with AI support (Assistant) and without (Control). Expert raters were blind to the condition and used the same scoring scale from our technical evaluation (Appendix A.1).